# An Efficient Sparse Kernel Generator for O(3)-Equivariant Deep Networks

### v1

@@ Abstract
Rotation equivariant graph neural networks, i.e. networks designed to guarantee certain geometric relations between their inputs and outputs, yield state of the art performance on spatial deep learning tasks. They exhibit high data efficiency during training and significantly reduced inference time for interatomic potential calculations compared to classical approaches. Key to these models is the Clebsch-Gordon (CG) tensor product, a kernel that contracts two dense feature vectors with a highly-structured sparse tensor to produce a dense output vector. The operation, which may be repeated millions of times for typical equivariant models, is a costly and inefficient bottleneck. We introduce a GPU sparse kernel generator for the CG tensor product that provides significant speedups over the best existing open and closed-source implementations. Our implementation achieves high performance by carefully managing the limited GPU shared memory through static analysis at model compile-time, minimizing reads and writes to global memory. We break the tensor product into a series of smaller kernels with operands that fit entirely into registers, enabling us to emit long arithmetic instruction streams that maximize instruction-level parallelism. By fusing the CG tensor product with a subsequent graph convolution , we reduce both intermediate storage and global memory traffic over naïve approaches that duplicate input data. We also provide optimized kernels for the gradient of the CG tensor product and a novel identity for the higher partial derivatives required to predict interatomic forces. Our fused kernels offer up to 4.5x speedup for the forward pass and 3x for the backward pass over NVIDIA’s closed-source cuEquivariance package, as well as $>10$x speedup over the widely-used e3nn package. We offer up to 5.3x inference-time speedup for the MACE chemistry foundation model over the original unoptimized version.
@@ 1 Introduction
Each nonzero block of $\mathcal{P}$ in Figure 1 is a structured tensor (see Figure 2 for illustrations of the nonzero pattern). Popular codes [9, 10] fill these blocks with explicit zeros and use optimized dense linear algebra primitives to execute the tensor product, performing unnecessary work in the process. By contrast, we use Just-in-Time (JIT) compilation to generate kernels that only perform work for nonzero tensor entries, achieving significantly higher throughput as block sparsity increases (see Table 2). While previous works [18, 16] use similar fine-grained approaches to optimize kernels for each block in isolation, we achieve high throughput by JIT-compiling a single kernel for the entire sparse tensor $\mathcal{P}$. By compiling kernels optimized for an entire sequence of nonzero blocks, we exhibit a degree of instruction level parallelism and data reuse that prior approaches do not provide.
@@ 2.2 Core Computational Challenge
Let $\bm{x}\in\mathbb{R}^{n},\bm{y}\in\mathbb{R}^{m}$ be two vectors from some intermediate layer $\bm{\phi}$ of an equivariant deep neural network. For example, $\bm{x}$ could be the embedding associated with a node of a graph and $\bm{y}$ a feature vector for an edge (see Figure 4, bottom right). We can view both vectors as functions $\bm{x}(\bm{v}),\bm{y}(\bm{v})$ of the network input $\bm{v}$, which are equivariant to $(\bm{D}_{\textrm{in}},\bm{D}_{x})$ and $(\bm{D}_{\textrm{in}},\bm{D}_{y})$ respectively. An equivariant graph convolution layer $\bm{\phi}$ interacts $\bm{x}$ and $\bm{y}$ to produce a new vector $\bm{z}$. To ensure layer-equivariance of $\phi$, $\bm{z}(\bm{v})$ must be equivariant to $(\bm{D}_{\textrm{in}},\bm{D}_{z})$, where $\bm{D}_{z}$ is a new representation selected by the network designer.
@@
Let $\bm{P}\in\mathbb{R}^{nm\times nm}$ be the similarity transform diagonalizing $\bm{D}_{x}\otimes\bm{D}_{y}$. To reduce the dimension of the Kronecker product, we first form $\bm{P}(\bm{x}(\bm{v})\otimes\bm{y}(\bm{v}))$, which is an equivariant function with output representation $\bm{P}^{-1}(\bm{D}_{x}\otimes\bm{D}_{y})\bm{P}$. We can now safely remove segments of $\bm{P}(\bm{x}\otimes\bm{y})$ corresponding to unneeded higher-order Wigner blocks and recombine its components through a trainable, structured weight matrix. The result, $\bm{z}(\bm{v})=\bm{W}\bm{P}(\bm{x}(\bm{v})\otimes\bm{y}(\bm{v}))$, has a new output representation $\bm{D}_{z}$ and can be much shorter than $\bm{x}\otimes\bm{y}$.
@@
Our goal is to accelerate computation of $\textrm{TP}(\mathcal{P},\bm{x},\bm{y},\bm{W})$ for a variety of CG coefficent tensors $\mathcal{P}$. Given $\partial E/\partial\bm{z}$ for some scalar quantity $E$, we will also provide an efficient kernel to compute $\partial E/\partial\bm{x}$, $\partial E/\partial\bm{y}$, and $\partial E/\partial\bm{W}$ in a single pass. These gradients are required during both training and inference for some interatomic potential models.
@@ 2.3 Structure in the Sparse Tensor and Weights
Instead, the CG tensor product splits into a sequence of subkernels [18] that match the structure in $\mathcal{P}$ and $\bm{W}$. We target two common patterns. First, the CG tensor product may interact $b$ unique segments of $\bm{x}$ with a common segment of $\bm{y}$ using the same CG sparse tensor, followed by multiplication by a submatrix of weights from $\bm{W}$ rearranged along a diagonal. Figure 3B illustrates this primitive as a contraction of a sparse tensor with a matrix $\bm{X}$ (containing the $b$ rearranged segments from $\bm{x}$) and the common vector $\bm{y}$. It appears in the Nequip [4] and MACE [2] models. Kernel C is identical to kernel B, but arranges the weights in a dense matrix $\bm{W}\in\mathbb{R}^{b^{\prime}\times b}$; the row count of $\bm{Z}$ may be distinct from the row count of $\bm{X}$. The latter operation appears in DiffDock [6] and 3D shape classifiers by Thomas et al. [32]. Both operations can be extended to interact multiple segments from $\bm{y}$, but we could not find this pattern in existing equivariant models. While kernels besides B and C are possible, they rarely appear in practice.
@@ 2.5 Context and Related Work
The cuEquivariance package [10], released by NVIDIA concurrent to this project’s development, offers the fastest implementation of the CG tensor product outside of our work. However, their kernels are closed-source and do not appear to exploit sparsity within each nonzero block of $\mathcal{P}$. Our kernel performance matches or exceeds cuEquivariance, often by substantial margins. Related codes include the GELib library [18] and Equitriton [22], the former to efficiently compute tensor products and the latter to evaluate spherical harmonic polynomials. Although GELib is less widely used in production models, we analyze its custom CUDA kernels in Section 3.6.
@@
The intense cost of the CG tensor product has fueled the search for cheaper yet accurate algorithms. For example, Schütt et al. [29] propose an equivariant message passing neural network that operates in Cartesian space, eliminating the need for CG tensor products. Luo et al. [25] connect the tensor product with a spherical harmonic product accelerated through the fast Fourier transform, an operation they call the Gaunt tensor product. Xie et al. [36] counter that the Gaunt tensor product does not produce results directly comparable to the CG tensor product, and that the former may sacrifice model expressivity. Lee et al. [21] observe that higher-order equivariant features may not improve model performance, while Qu and Krishnapriyan [28] propose eliminating equivariance in interatomic potential models altogether.
@@ 3 Engineering Efficient CG Kernels
Our task is to generate an efficient CG tensor product kernel given a problem specification outlined in Section 2.4. Algorithm 1 describes the logic of a kernel that operates on a large batch of inputs, each with a distinct set of weights (see Figure 4B). We assign each $(\bm{x},\bm{y},\bm{W})$ input triple to a single GPU warp, a choice which has two consequences. First, it enables each warp to execute contiguous global memory reads / writes for $\bm{x},\bm{y},\bm{W}$ and $\bm{z}$. Second, it allows warps to execute in a completely asynchronous fashion without any CTA-level synchronization, boosting throughput significantly. The weights $\bm{W}$ are stored in a compressed form without the zero entries illustrated in Figure 1.
After the division of labor, each warp follows a standard GPU kernel workflow. The three inputs are staged in shared memory, the kernels in Equation (2.2) are executed sequentially, and each output $\bm{z}_{b}$ is stored back. Each warp operates on a unique partition of the CTA shared memory which may not be large enough to contain the the inputs and outputs. In the latter case, chunks of $\bm{x}_{b}$, $\bm{y}_{b}$, $\bm{W}_{b}$, and $\bm{z}_{b}$ are staged, and the computation executes in phases according to a schedule described in Section 3.1.
@@ 3.2 JIT Subkernel Implementations
In preperation to execute a subkernel, suppose we have loaded $\bm{x}$, $\bm{y}$ and $\bm{W}$ into shared memory and reshaped subranges of all three to form $\bm{X}_{\textrm{kern}},\bm{y}_{\textrm{kern}}$, and $\bm{W}_{\textrm{kern}}$. Algorithm 2 gives the pseudocode to execute either kernel B or C from Figure 3 using these staged operands.
@@
The matrix multiplication by the weights at the end of Algorithm 2 depends on the structure of $\bm{W}$. When $\bm{W}$ is square adn diagonal (kernel B), multiplication proceeds asynchronously in parallel across all threads. When $\bm{W}$ is a general dense matrix, we temporarily store $\bm{z}_{\textrm{reg}}$ to shared memory and perform a warp-level matrix-multiplication across all threads. Our final implementation uses a mix of hand-rolled kernels and CUTLASS [31] for warp matrix multiplication.
@@ 3.4 Higher Partial Derivatives
Now our task is to compute ($\partial\mathcal{L}/\partial\bm{x}$, $\partial\mathcal{L}/\partial\bm{y}$, $\partial\mathcal{L}/\partial\bm{W}$, $\partial\mathcal{L}/\partial\bm{d}$) given ($\partial\mathcal{L}/\partial\bm{a}$, $\partial\mathcal{L}/\partial\bm{b}$, $\partial\mathcal{L}/\partial\bm{C}$). We dispatch seven calls to the forward and backward pass kernels:
(3.4) op1 $=\textrm{backward}(\bm{a},\bm{b},\bm{W},\bm{g}_{z})$
op2 $=\textrm{backward}(\bm{x},\bm{y},\bm{C},\bm{g}_{z})$
op3 $=\textrm{TP}(\bm{a},\bm{y},\bm{W})$
op4 $=\textrm{backward}(\bm{a},\bm{y},\bm{W},\bm{g}_{z})$
op5 $=\textrm{backward}(\bm{x},\bm{b},\bm{W},\bm{g}_{z})$
op6 $=\textrm{TP}(\bm{x},\bm{b},\bm{W})$
op7 $=\textrm{TP}(\bm{x},\bm{y},\bm{C}).$
By repeatedly applying the product and chain rules to the formulas for $a$, $b$, and $c$ in Section 3.3, we can show
@@ 4 Experiments
Our kernel generator is available online [^1] https://github.com/vbharadwaj-bk/OpenEquivariance. We adopted the frontend interface of e3nn [8, 9] and used QuTiP [13, 20] to generate CG coefficients; we thank the authors of both packages. We tested correctness against e3nn to ensure that our kernels produce identical results, up to floating point roundoff and a well-defined reordering of the weights $\bm{W}$ on certain input configurations.
@@
As baselines, we used the PyTorch versions of e3nn (v0.5.4) [9] and NVIDIA cuEquivariance (v0.0.1) [10]. A documented issue in the cuEquivariance kernel package prevented us from benchmarking their backward pass for most configurations excluding MACE. We will update our benchmarks when these issues are resolved. The e3nn implementation was accelerated with torch.compile except where prohibited by memory constraints.
In Figures 5 and 6, we benchmarked all functions through a uniform PyTorch interface and included any overhead in the measured runtime. Figures 7, Table 2, and Figure 9 (right) rely on kernel runtime measurements without Pytorch overhead.
@@ 4.1 Throughput Comparison
Figure 5 shows the results of our profiling on configurations that use only Kernel B (see Figure 3). Our median FP32 speedup over e3nn was 6.8x, resp. 4.7x, for the forward and backward passes, with a maximum speedup of 10.2x for the forward pass. We observed a median speedup of 1.7x over cuE for the FP32 forward pass, which drops to 1.3x in FP64 precision. We achieve parity with cuE on the larger MACE model in FP64 precision; here, both cuE and our kernels have nearly identical runtime. We explore this configuration further in our kernel fusion benchmarks. Our implementation exhibits an outsized speedup (10-20x) on the Nequip-benzene configuration over both baselines. Here, the large output vector size ($\sim 20\textrm{k}$ data words) requires careful shared memory management, which our scheduling algorithm provides.
To benchmark kernel C, we used the Tetris polynomial from e3nn’s documentation [9] and two configurations based on DiffDock [6]. We exhibit between 1.5x and 2.1x speedup over cuE on the forward and backward passes on DiffDock. The Tetris polynomial uses an identical weight matrix $\bm{W}$ for all inputs $\bm{x}_{1}...\bm{x}_{B}$ and $\bm{y}_{1}...\bm{y}_{B}$, a feature unique to this model. We use atomic accumulation in the backward pass to compute the gradient with respect to the shared weights, which reduces our relative speedup in FP64 precision. Our speedups over all configurations are less dramatic for kernel C, which has a workload dominated by the small dense matrix multiplication in Algorithms 2 and 3.

### v3

@@ Abstract
Rotation equivariant graph neural networks, i.e. networks designed to guarantee certain geometric relations between their inputs and outputs, yield state of the art performance on spatial deep learning tasks. They exhibit high data efficiency during training and significantly reduced inference time for interatomic potential calculations compared to classical approaches. Key to these models is the Clebsch-Gordon (CG) tensor product, a kernel that contracts two dense feature vectors with a highly-structured sparse tensor to produce a dense output vector. The operation, which may be repeated millions of times for typical equivariant models, is a costly and inefficient bottleneck. We introduce a GPU sparse kernel generator for the CG tensor product that provides significant speedups over the best existing open and closed-source implementations. Our implementation achieves high performance by carefully managing the limited GPU shared memory through static analysis at model compile-time, minimizing reads and writes to global memory. We break the tensor product into a series of smaller kernels with operands that fit entirely into registers, enabling us to emit long arithmetic instruction streams that maximize instruction-level parallelism. By fusing the CG tensor product with a subsequent graph convolution , we reduce both intermediate storage and global memory traffic over naïve approaches that duplicate input data. We also provide optimized kernels for the gradient of the CG tensor product and a novel identity for the higher partial derivatives required to predict interatomic forces. Our fused kernels offer up to 4.8x speedup for the forward pass and 3x for the backward pass over NVIDIA’s closed-source cuEquivariance package, as well as $>10$x speedup over the widely-used e3nn package. We offer up to 5.4x inference-time speedup for the MACE chemistry foundation model over the original unoptimized version.
@@ 1 Introduction
Each nonzero block of $\mathcal{P}$ in Figure 1 is a structured sparse tensor (see Figure 2 for illustrations of the nonzero pattern). Popular codes [10, 11] fill these blocks with explicit zeros and use optimized dense linear algebra primitives to execute the tensor product, performing unnecessary work in the process. By contrast, we use Just-in-Time (JIT) compilation to generate kernels that only perform work for nonzero tensor entries, achieving significantly higher throughput as block sparsity increases (see Table 2). While previous works [19, 17] use similar fine-grained approaches to optimize kernels for each block in isolation, we achieve high throughput by JIT-compiling a single kernel for the entire sparse tensor $\mathcal{P}$. By compiling kernels optimized for an entire sequence of nonzero blocks, we exhibit a degree of instruction level parallelism and data reuse that prior approaches do not provide.
@@ 2.2 Core Computational Challenge
Let $\bm{x}\in\mathbb{R}^{n},\bm{y}\in\mathbb{R}^{m}$ be two vectors from some intermediate layer $\bm{\phi}$ of an equivariant deep neural network. For example, $\bm{x}$ could be the embedding associated with a node of a graph and $\bm{y}$ a feature vector for an edge (see Figure 4, bottom). We can view both vectors as functions $\bm{x}(\bm{v}),\bm{y}(\bm{v})$ of the network input $\bm{v}$, which are equivariant to $(\bm{D}_{\textrm{in}},\bm{D}_{x})$ and $(\bm{D}_{\textrm{in}},\bm{D}_{y})$ respectively. An equivariant graph convolution layer $\bm{\phi}$ interacts $\bm{x}$ and $\bm{y}$ to produce a new vector $\bm{z}$. To ensure layer-equivariance of $\phi$, $\bm{z}(\bm{v})$ must be equivariant to $(\bm{D}_{\textrm{in}},\bm{D}_{z})$, where $\bm{D}_{z}$ is a new representation selected by the network designer.
@@
Let $\bm{P}\in\mathbb{R}^{nm\times nm}$ be the similarity transform diagonalizing $\bm{D}_{x}\otimes\bm{D}_{y}$. To reduce the dimension of the Kronecker product, we first form $\bm{P}(\bm{x}(\bm{v})\otimes\bm{y}(\bm{v}))$, an equivariant function with block-diagonal output representation $\bm{P}(\bm{D}_{x}\otimes\bm{D}_{y})\bm{P}^{-1}$. We can now safely remove segments of $\bm{P}(\bm{x}\otimes\bm{y})$ corresponding to unneeded higher-order Wigner blocks and recombine its components through a trainable, structured weight matrix. The result, $\bm{z}(\bm{v})=\bm{W}\bm{P}(\bm{x}(\bm{v})\otimes\bm{y}(\bm{v}))$, has a new output representation $\bm{D}_{z}$ and can be much shorter than $\bm{x}\otimes\bm{y}$.
@@
Our goal is to accelerate computation of $\textrm{TP}(\mathcal{P},\bm{x},\bm{y},\bm{W})$ for a variety of CG coefficient tensors $\mathcal{P}$. Given $\partial E/\partial\bm{z}$ for some scalar quantity $E$, we will also provide an efficient kernel to compute $\partial E/\partial\bm{x}$, $\partial E/\partial\bm{y}$, and $\partial E/\partial\bm{W}$ in a single pass. These gradients are required during both training and inference for some interatomic potential models.
@@ 2.3 Structure in the Sparse Tensor and Weights
Instead, the CG tensor product splits into a sequence of subkernels [19] that match the structure in $\mathcal{P}$ and $\bm{W}$. We target two common patterns. First, the CG tensor product may interact $b$ unique segments of $\bm{x}$ with a common segment of $\bm{y}$ using the same block from $\mathcal{P}$, followed by multiplication by a submatrix of weights from $\bm{W}$ rearranged along a diagonal. Figure 3B illustrates Kernel B as a contraction of a sparse tensor with a matrix $\bm{X}$ (containing the $b$ rearranged segments from $\bm{x}$) and the common vector $\bm{y}$. It appears in the Nequip [4] and MACE [2] models. Kernel C is identical to kernel B, but arranges the weights in a dense matrix $\bm{W}\in\mathbb{R}^{b^{\prime}\times b}$. Here, $\bm{Z}$ and $\bm{X}$ may have distinct row counts. The latter operation appears in DiffDock [7] and 3D shape classifiers by Thomas et al. [34]. Both operations can be extended to interact multiple segments from $\bm{y}$, but we could not find this pattern in existing equivariant models. While kernels besides B and C are possible, they rarely appear in practice.
@@ 2.5 Context and Related Work
The cuEquivariance package [11], released by NVIDIA concurrent to this project’s development, offers the fastest implementation of the CG tensor product outside of our work. However, their kernels are closed-source and do not appear to exploit sparsity within each nonzero block of $\mathcal{P}$. Our kernel performance matches or exceeds cuEquivariance, often by substantial margins. Other relevant codes include GELib [19], Sphericart [5], and Equitriton [23]. The former efficiently computes CG tensor products, while the latter two accelerate spherical harmonic polynomial evaluation. Although GELib is less widely used in production models, we analyze its custom CUDA kernels in Section 3.6.
@@
The intense cost of the CG tensor product has fueled the search for cheap yet accurate algorithms. For example, Schütt et al. [31] propose an equivariant message passing neural network that operates in Cartesian space, eliminating the need for CG tensor products. For SO(3)-equivariant graph convolution, Passaro and Zitnick [28] align the node embeddings with spherical harmonic edge features before interacting the two. This innovation sparsifies the interaction tensor and asymptotically decreases the cost of each tensor product. Luo et al. [26] also produce asymptotic speedups by connecting the tensor product with a spherical harmonic product accelerated through the fast Fourier transform, an operation they call the Gaunt tensor product. Xie et al. [38] counter that the Gaunt tensor product does not produce results directly comparable to the CG tensor product, and that the former may sacrifice model expressivity. Lee et al. [22] observe that higher-order equivariant features may not improve model performance, while Qu and Krishnapriyan [30] achieve high accuracy by eliminating hard rotational constraints in atomic potential models.
@@ 3 Engineering Efficient CG Kernels
Our task is to generate an efficient CG tensor product kernel given a problem specification outlined in Section 2.4. Algorithm 1 describes the logic of a kernel that operates on a large batch of inputs, each with a distinct set of weights (see Figure 4A). We assign each $(\bm{x},\bm{y},\bm{W})$ input triple to a single GPU warp, a choice which has two consequences. First, it enables each warp to execute contiguous global memory reads / writes for $\bm{x},\bm{y},\bm{W}$ and $\bm{z}$. Second, it allows warps to execute in a completely asynchronous fashion without any CTA-level synchronization, boosting throughput significantly. The weights $\bm{W}$ are stored in a compressed form without the zero entries illustrated in Figure 1.
After the division of labor, each warp follows a standard GPU kernel workflow. The three inputs are staged in shared memory, the kernels in Equation (2.2) are executed sequentially, and each output $\bm{z}_{b}$ is stored back. Each warp operates on a unique partition of the CTA shared memory which may not be large enough to contain the the inputs and outputs. In the latter case, chunks of $\bm{x}$, $\bm{y}$, $\bm{W}$, and $\bm{z}$ are staged, and the computation executes in phases according to a schedule described in Section 3.1.
@@ 3.2 JIT Subkernel Implementations
In preparation to execute a subkernel, suppose we have loaded $\bm{x}$, $\bm{y}$ and $\bm{W}$ into shared memory and reshaped subranges of all three to form $\bm{X}_{\textrm{kern}},\bm{y}_{\textrm{kern}}$, and $\bm{W}_{\textrm{kern}}$. For the rest of Section 3.2 and all of Section 3.3, we omit the “kern” subscripts. Algorithm 2 gives the pseudocode to execute either kernel B or C from Figure 3 using these staged operands.
@@
The matrix multiplication by the weights at the end of Algorithm 2 depends on the structure of $\bm{W}$. When $\bm{W}$ is square and diagonal (kernel B), multiplication proceeds asynchronously in parallel across all threads. When $\bm{W}$ is a general dense matrix, we temporarily store $\bm{z}_{\textrm{reg}}$ to shared memory and perform a warp-level matrix-multiplication across all threads. Our final implementation uses a mix of hand-rolled kernels and CUTLASS [33] for warp matrix multiplication.
@@ 3.4 Higher Partial Derivatives
Now our task is to compute ($\partial\mathcal{L}/\partial\bm{x}$, $\partial\mathcal{L}/\partial\bm{y}$, $\partial\mathcal{L}/\partial\bm{W}$, $\partial\mathcal{L}/\partial\bm{g}_{z}$) given ($\partial\mathcal{L}/\partial\bm{a}$, $\partial\mathcal{L}/\partial\bm{b}$, $\partial\mathcal{L}/\partial\bm{C}$). We dispatch seven calls to the forward and backward pass kernels:
(3.4) op1 $=\textrm{backward}(\partial\mathcal{L}/\partial\bm{a},\partial\mathcal{L}/\partial\bm{b},\bm{W},\bm{g}_{z})$
op2 $=\textrm{backward}(\bm{x},\bm{y},\partial\mathcal{L}/\partial\bm{C},\bm{g}_{z})$
op3 $=\textrm{TP}(\partial\mathcal{L}/\partial\bm{a},\bm{y},\bm{W})$
op4 $=\textrm{backward}(\partial\mathcal{L}/\partial\bm{a},\bm{y},\bm{W},\bm{g}_{z})$
op5 $=\textrm{backward}(\bm{x},\partial\mathcal{L}/\partial\bm{b},\bm{W},\bm{g}_{z})$
op6 $=\textrm{TP}(\bm{x},\partial\mathcal{L}/\partial\bm{b},\bm{W})$
op7 $=\textrm{TP}(\bm{x},\bm{y},\partial\mathcal{L}/\partial\bm{C}).$
By repeatedly applying the product and chain rules to the formulas for $\bm{a}$, $\bm{b}$, and $\bm{C}$ in Section 3.3, we can show
@@ 4 Experiments
Our kernel generator is available online111 https://github.com/PASSIONLab/OpenEquivariance as an installable Python package. We adopted the frontend interface of e3nn [9, 10] and used QuTiP [14, 21] to generate CG coefficients; we thank the authors of both packages. We tested correctness against e3nn to ensure that our kernels produce identical results, up to floating point roundoff and a well-defined permutation of the weights $\bm{W}$ on certain input configurations. In cases where weight reordering is required, we provide a function for easy migration. We use the NVIDIA Runtime Compiler (NVRTC) to compile our generated kernels through a C++ extension to Python.
@@
As baselines, we used the PyTorch versions of e3nn (v0.5.4) [10] and NVIDIA cuEquivariance (v0.0.2) [11]. The e3nn implementation was accelerated with torch.compile except where prohibited by memory constraints. For Figures 5 and 6, we benchmarked all functions through a uniform PyTorch interface and included any overhead in the measured runtime. Figures 7, Table 2, and Figure 9 (right) rely on kernel runtime measurements without PyTorch overhead.
@@ 4.1 Throughput Comparison
Figure 5 shows the results of our profiling on configurations that use only Kernel B (see Figure 3). Our median FP32 speedup over e3nn was 5.7x, resp. 5.0x, for the forward and backward passes, with a maximum of 9.3x for the forward pass. We observed a median speedup of 1.7x over cuE for the FP32 forward pass, which drops to 1.2x in FP64 precision. We achieve parity with cuE on the larger MACE model in FP64 precision; here, both cuE and our kernels have nearly identical runtime. We explore this configuration further in our kernel fusion benchmarks. Our implementation exhibits an outsized speedup (10-20x) on the Nequip-benzene configuration over both baselines. Here, the large output vector size ($\sim 20\textrm{k}$ data words) requires careful shared memory management, which our scheduling algorithm provides.
To benchmark kernel C, we used the Tetris polynomial from e3nn’s documentation [10] and two configurations based on DiffDock [7]. We exhibit between 1.4x and 2.1x speedup over cuE for both forward and backward passes on DiffDock. The Tetris polynomial uses an identical weight matrix $\bm{W}$ for all inputs $\bm{x}_{1}...\bm{x}_{B}$ and $\bm{y}_{1}...\bm{y}_{B}$, a feature unique to this model. We use atomic accumulation in the backward pass to compute the gradient with respect to the shared weights, which reduces our relative speedup in FP64 precision. Our speedups over all configurations are less dramatic for kernel C, which has a workload dominated by the small dense matrix multiplication in Algorithms 2 and 3.